OpenStudy (anonymous):

Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid $130. This year, he paid $143. What was the percent increase in the cost of his plane ticket? Show your work.

OpenStudy (anonymous):

increase was 10%

OpenStudy (anonymous):

change in price is 143-130 = 13 13/130 x 100% = 10%

OpenStudy (anonymous):

check 130 plus 10% of 130 = 130 + 13 = 143